OpenStudy (swimmerranae102000):

Can someone walk me through this please? (x+2)+(2x-1)=4x-7

OpenStudy (kewlgeek555):

I'm so sorry if I am keeping you waiting. I am trying to work this out to see if I wil get the answer right so I can correctly walk you through it. I haven't done math in a while. :P

OpenStudy (swimmerranae102000):

Oh its ok! Thanks

OpenStudy (joshtray):

Open the bracket first x+2x=3x and 2-1=1, So we have\[3x+1=4x-7\]

OpenStudy (swimmerranae102000):

i dont understand how you got that tbh...

OpenStudy (joshtray):

Collecting like terms, we have\[4x-3x=7+1\]

OpenStudy (joshtray):

What is tbh?

OpenStudy (swimmerranae102000):

to be honest sorry

OpenStudy (swimmerranae102000):

ok so how do i find like terms...like i need baby steps on this

OpenStudy (joshtray):

\[(x+2)+(2x-1)=4x-7\] \[x+2+2x-1=4x-7\] \[x+2x+2-1=4x-7\]

OpenStudy (swimmerranae102000):

so you flipped x+2 and x+2x around? IM RETARDED IM SO SORRY RN!

OpenStudy (kewlgeek555):

Oh wait you need to find x, don't you? Okay. Well first we would have to simplify both sides of the equation. So let's start with (x + 2) + (2x + 1). You would have to use what I believe is the distributive property to simplify this. You use imaginary 1s to take out the parenthesis. 1 * (x + 2) + 1 * (2x + 1) x + 2 + 2x -1 Now that you have taken away the parenthesis, you combine like terms, 3x + 1 Now the equation is 3x + 1 = 4x - 7. Now you will have to isolate the variable. You will do this by putting the variable on one side and the numbers on the other. 3x + 1 = 4x - 7 I can do this part with you since it's a little bit more difficult, would you know maybe how to isolate the variable?

OpenStudy (joshtray):

Something like that, yeah

OpenStudy (swimmerranae102000):

still a little confused..for real im sorry if i seem retarded but these are one of the things i dont get

OpenStudy (mathmale):

Please, when you post something like "(x+2)+(2x-1)=4x-7," would you include the instructions? You yourself might understand the question better if you were clear about your goal here. Removing parentheses, x + 2 + 2x - 1 = 4x - 7 combine like terms on the left side. Your result, please?

OpenStudy (swimmerranae102000):

idk what like terms are..idk how to get them...

OpenStudy (swimmerranae102000):

Given that A is between B and C. If AB = x+2, BC = 4x-7, and AC = 2x -1, find the x and AC. Could you explain because im having trouble? This is the original question but i plugged the numbers in and that part idk how to get

OpenStudy (mathmale):

kewlgeek555 has already offered help with this: 3x + 1 = 4x - 7 x and 2x are "like" terms, as are 2 and -1. What is the sum of x and 2x? What is the sum of 2 and -1? If you do this correctly you will get the same result as kewlgeek555 has already gotten.

OpenStudy (kewlgeek555):

@swimmerranae102000 , @Joshtray actually used the distributive property, just like I did, and took out the parenthesis with the imaginary 1s. When you have something like (x+2)+(2x-1), you should know that there are imaginary 1s that you can distribute (by multiplying) to take away the parenthesis.